\begin{thebibliography}{99}
\bibitem{Bertsimas}
  D. Bertsimas, I. Popescu, {\it Optimal inequalities in probability
theory: a convex optimization approach,}  SIAM J. Optim.  {\bf 15}
(2005),  no. 3, 780--804.

\bibitem{HP}J.W. Helton, M. Putinar,  {\it  Positive Polynomials in Scalar and
Matrix Variables, the Spectral Theorem and Optimization}, in vol.
{\it Operator Theory, Structured Matrices, and Dilations}, Theta,
Bucharest, 2007, pp 229-306.


\bibitem{Kakutani}
S. Kakutani, {\it Ein Beweis des Satzes von M. Eidelheit \"uber
konvexe Mengen,} Proc. Imp. Acad. Tokyo {\bf 13}(1937), 93-94.


  \bibitem{Ko} G. K\"othe,  \textit {Topological Vector Spaces.
  I.} Springer, Berlin, 1969.

  \bibitem{KR} Krein, M.G., Rutman, M.A., {\it Linear operators
   leaving invariant a cone in a Banach space} (in Russian) Uspehi
   Mat. Nauk {\bf 23}(1948), pp. 3-95.



\bibitem{Lasserre0} J. B. Lasserre,
{\it Optimisation globale et th‰orie des moments}, C. R. Acad.
Sci. Paris {\bf 331} S\'erie 1, 929--934.


\bibitem{Lasserre1}
J.B. Lasserre, {\it Global optimization with polynomials and the
problem of moments},  SIAM J. Optim. {\bf 11}(2001), 796-817.

\bibitem{LP} J. -B. Lasserre, M. Putinar, {\it Positivity and optimization
for semi-algebraic functions}, arXiv:0910.5250 

 \bibitem{posbook} {\it Positive Polynomials in Control}, (D.
 Henrion, A. Garulli, eds.),
Lecture Notes in Control and Information Sciences, Springer,
Berlin, Heidelberg, 2005.

  \bibitem{PD} A. Prestel, C. Delzell, {\it Positive polynomials},
  Springer, Berlin, 2001.

   \bibitem{Put} M. Putinar,
   {\it Positive polynomials on compact semi-algebraic sets}, Indiana
   Univ. Math. J. {\bf 42}, (1993) 969-984
   

   \bibitem{alggeom} {\it Emerging Applications of Algebraic Geometry},
   (M. Putinar, S. Sullivant, eds.),
IMA Series in Applied Mathematics, Springer, Berlin, 2008.



\bibitem{RN} F. Riesz, B. Sz.-Nagy,
{\it Functional analysis},  Transl. from the 2nd French ed. by Leo
F. Boron. Reprint of the 1955 orig. publ. by Ungar Publ. Co.,
Dover Books on Advanced Mathematics, Dover Publications, Inc., New
York, 1990.

\bibitem{MRiesz}
M. Riesz, {\it  Sur le probl\`eme des moments. Troisi\`eme Note},
Ark.\ Mat.\ Fys.\ \textbf{16}(1923), 1--52.

  \bibitem{Sc} K. Schm\"udgen,
  {\it The K-moment problem for compact semi-algebraic sets}, Math. Ann.
{\bf 289} (1991) 203-206.

  \end{thebibliography}

\documentstyle[11pt]{article}

\newtheorem{theorem}{Theorem}[section]

\newtheorem{e-proposition}[theorem]{Proposition}
\newtheorem{corollary}[theorem]{Corollary}
\newtheorem{e-definition}[theorem]{Definition\rm}

\begin{document}

 \begin{center}{\bf A Striktpositivstellensatz for measurable functions 
 
 (corrected version)}\bigskip

{Mihai Putinar}\bigskip

{Mathematics Department, 

University of California,

Santa Barbara, CA 93106,

{\rm mputinar@math.ucsb.edu}}
\end{center}

\begin{abstract}
A weighted sums of squares decomposition
of positive Borel measurable functions on a bounded Borel subset
of the Euclidean space is obtained via duality from the spectral
theorem for tuples of commuting self-adjoint operators. The
analogous result for polynomials or certain rational functions was
amply exploited during the last decade in a variety of
applications.

\end{abstract}

\section{Introduction}
To put our main result into the current real algebra context, we
recall below the abstract framework for studying linear
decompositions into weighted sums of squares.

Let $A$ be a commutative algebra with 1, over the rational field.
A {\it quadratic module} $Q \subset A$ is a subset of $A$ such
that $Q+Q \subset Q, 1\in Q$ and $a^2Q\subset A$ for all $a\in A$.
We denote by $Q(F; A)$ or simply $Q(F)$ the quadratic module
generated in $A$ by the set $F$. That is $Q(F; A)$ is the smallest
subset of $A$ which is closed under addition and multiplication by
squares $a^2, \ a \in A$, containing $M$ and the unit $1 \in A$.
If $F$ is finite, we say that the quadratic module is finitely
generated. A quadratic module which is also closed under
multiplication is called a {\it quadratic preordering}. A
quadratic module $Q$ is called archimedean if the constant
function $1$ belong to its algebraic interior, that is, for every
$f \in Q$ there exists $\epsilon >0$ such that $1+tf \in Q$ for
all $0 \leq t \leq \epsilon.$

Assume that $A = {{\mathbf R}}[x_1,...,x_d]$ is the polynomial
algebra. The positivity set $P(Q)$ of $Q \subset {\mathbf
R}[x_1,...,x_d] $ is the set of all points $x \in {\mathbf{R}}^d$
for which $q(x) \geq 0, \  q \in Q$.

The following Striktpositivestellensatz has attracted in the last
decade a lot of attention from practitioners of polynomial
optimization:
{\it Let $Q\subset {{\mathbf R}}[x_1,...,x_d]$ be an
archimedean quadratic module and assume that a
polynomial $f$ is positive on $P(Q)$. Then $f \in Q$.}

This fact was discovered by the author \cite{Put}, generalizing
Schm\"udgen's Striktpositivestellensatz \cite{Sc} for the {\it
finitely generated preordering} associated to a compact
non-negativity set.

In plain language, the above result can be stated as follows.
Denote by $\Sigma^2$ the convex cone of sums of squares in the
polynomials ring ${{\mathbf R}}[x_1,...,x_d]$. Let $p_1,...,p_k \in
{{\mathbf R}}[x_1,...,x_d]$ be polynomials, so that the quadratic
module generated by them $Q(p_1,...,p_k) = \Sigma^2 + p_1 \Sigma^2
+...+ p_k \Sigma^2$ is archimedean. That is, there exists
$\epsilon>0$ so that $ 1-\epsilon (x_1^2+...+x_d^2) \in Q,$ (for
the reduction of $Q$ archimedean to this criterion see \cite{PD}).
The stated Striktpositivstellensatz asserts: if a polynomial $P$
is positive on the set $ S(p_1,...,p_k) =\{ x = (x_1,...,x_d); \
p_i(x) \geq 0, \  1 \leq i \leq k\},$ then $P \in Q(p_1,...,p_k)$.

A simple duality argument implies under the above conditions that
every linear functional $L \in {\mathbf{R}}[x]'$ which is
non-negative on the quadratic module $Q(p_1,...,p_k)$ is
represented by a positive Borel measure $\mu$, supported by the
basic semi-algebraic set $S(p_1,...,p_k)$:
$$ L(f) = \int_{S(p_1,...,p_k)} f d\mu, \ \  f \in
{\mathbf{R}}[x].$$ In general, a quadratic module $Q \subset A$,
where $A$ is an algebra of measurable functions defined on 
${\mathbf R}^d$ has {\it the moment property} if every linear functional
on $A$ which is non-negative on $q$ is represented by a positive measure
supported by $P(Q)$.

The correspondence between the above
Positivestellensatz and the multivariate moment problem with
prescribed compact semi-algebraic supports works fruitfully in
both directions. First, the original proof of the
Positivstellensatz was obtained via the standard moment problem
solution offered by the spectral theorem (see \cite{Put} and
\cite{PD} for an algebraic proof). Second, and more important for
applications, it was J. B. Lasserre \cite{Lasserre0,Lasserre1} who
has interpreted the moments
$$y_\alpha = \int x^\alpha d\mu, \ \  \alpha \in \mathbf{N}^d,$$
of the representing measure as new independent variables and has
obtained a hierarchy of linear, semi-definite optimization problem
converging to the minimization of a given polynomial on a compact
semi-algebraic set. For more details towards applications and
theoretical ramifications see \cite{Bertsimas,HP,posbook,alggeom}.

\section{Main result} The aim of this note is to prove a
natural extension of the polynomial Positivstellensatz to algebras
of Borel measurable functions defined on Euclidean space. Although
a more general statement, on an arbitrary locally compact space or
even on a non-commutative $C^\ast$-algebra is possible to deduce
with similar techniques, we consider that the Euclidean space
setting is: first, the most important for applications, and
second, it contains a specific feature which makes it worth a
separate discussion.

\begin{theorem} Let $Q$ be a countably generated archimedean quadratic module
contained in the algebra \\
${\mathcal{A}} = {\mathbf{R}}[x_1,...,x_d,h_1,...,h_m]$ spanned by the
coordinate functions and by Borel measurable functions
$h_1,...,h_m$ on ${\mathbf{R}}^d$. Assume that $Q$ possesses the moment property. 
If a function $f \in {\mathcal{A}}$
is positive on $P(Q)$, then $f \in Q$.
\end{theorem}

{\bf Proof.} Since $Q$ is archimedean, there exists $\epsilon>0$
such that $1 -\epsilon(x_1^2+...+x_d^2+h_1^2+...+h_m^2) \in Q$.
Thus the positivity set $P(Q)$ is contained in the ball
$x_1^2+...+x_d^2 \leq 1/\epsilon.$ Because $Q$ is countably
generated, the set $P(Q)$ is Borel measurable.

The fact that $Q$ is archimedean as a convex cone, means that for
every $h \in {\mathcal{A}}$ there exists positive constants $c,C$
with the property $C-h, h-c \in Q$. Assume by contradiction that
the function $f$ does not belong to $Q$. According to Marcel Riesz
extension of positive functionals \cite{MRiesz}(known and
rediscovered over the years by many authors
\cite{Kakutani,KR,Ko}), there exists $L \in {\mathcal{A}}'$ so that:
$$ L(f) \leq 0 \leq L(q), \ L(1)>0, \ q \in Q.$$

Next we use Gelfand-Naimark-Segal construction of a Hilbert space
realization of the functional $L$. Specifically, $L(g^2) \geq 0$
for all $g \in {\mathcal{A}}$, and Cauchy-Schwarz' inequality proves
that the set $\mathcal{N}$ of functions $h \in {\mathcal{A}},
L(h)=0,$ is an ideal; whence we can introduce on the quotient
algebra ${\mathcal{A}}/\mathcal{N}$ the positive definite inner
product
$$ \langle g_1, g_2 \rangle = L(g_1 g_2),\ \ g_1,g_2 \in {\mathcal{A}}/\mathcal{N}.$$
Let $\mathcal{H}$ be the Hilbert space completion of
${\mathcal{A}}/\mathcal{N} \otimes_{{\mathbf R}} \mathbf{C}.$ Since
$Q$ is archimedean, the multiplication operators by each generator
$x_1,...,x_d, h_1,...,h_m$ extends by linearity to a tuple of
commuting {\it bounded} self-adjoint operators on $\mathcal{H}$,
denoted by $(X,H) = (X_1,...,X_d,H_1,...,H_m)$, respectively. In
view of the Spectral Theorem \cite{RN}, there exists a positive
measure $\sigma$ on ${\mathbf{R}}^{d+k}$, so that, for all bounded
Borel functions $F(x_1,...,x_d,y_1,...,y_m)$ we have
$$ \langle F(X_1,...,X_d,H_1,...,H_m)1, 1 \rangle= \int_{{\mathbf{R}}^{d+k}} F
d\sigma.$$ From here we deduce that $H_j = h_j(X_1,...,X_d), 1
\leq j \leq m,$ in the sense of Borel functional calculus for
self-adjoint operators \cite{RN}. Hence the measure $\sigma$ is
the push forward of a positive measure on ${\mathbf{R}}^d$ by the
graph map $ x \mapsto (x,h_1(x),...,h_m(x)), \ x \in
{\mathbf{R}}^d$:
$$ \langle F(X_1,...,X_d,H_1,...,H_m)1, 1 \rangle =
\int_{{\mathbf{R}}^{d}} F(x,h_1(x),...,h_m(x))d\mu(x).$$ This shows
that the measure $\mu$ has compact support, contained in the ball
centered at $0$, of radius $1/\epsilon$. Since the algebra $\mathcal A$ was supposed to possess the
moment property, the support of the measure $\mu$ is contained in the positivity set $P(Q)$.

Finally, recall from the statement that $f|_{P(Q)}>0$. On the
other hand, by the construction of the functional we have
$$ \int_{P(Q)} f d\mu = \int f d\mu=  \langle f(X,H) 1, 1 \rangle = L(f) \leq 0.$$
But $\mu({\mathbf{R}}^d)>0$, and thus we reach a contradiction.\\

The reader encounter no complications in specializing the theorem
above and its proof to a finitely generated quadratic module. We
simply state the result.

\begin{corollary}Let $q_1,...,q_n$ be elements of the algebra
 ${{\mathcal{A}}} = {\mathbf{R}}[x_1,...,x_d,h_1,...,h_m]$ generated by
the coordinate functions and by Borel measurable functions
$h_1,...,h_m$ on ${\mathbf{R}}^d$. Let $\Sigma {\mathcal{A}}^2$ denote
the convex cone of sums of squares, and consider the Borel
measurable set
$$ P(q_0, q_1,...,q_n) = \{ x \in {\mathbf{R}}^d; \ q_i(x) \geq 0, \
0
\leq i \leq n\},$$ where $q_0(x) = 1 -
(x_1^2+...+x_d^2+h_1^2+...+h_m^2)$.

If the quadratic module generated by $q_0,...,q_n$ has the moment property, then
a function $f \in {\mathcal{A}}$ is positive on $P(q_0,
q_1,...,q_n)$ satisfies $f \in \Sigma {\mathcal{A}}^2 + q_0 \Sigma
{\mathcal{A}}^2+...+q_n \Sigma {\mathcal{A}}^2.$\\
\end{corollary}

It is very natural to try to extend Lasserre's linearization and relaxation procedure to this
new framework. Specifically that means to consider the mixed moments
$$ y_{\alpha,\beta} = \int x^\alpha h(x)^\beta d\mu, \ \ \alpha
\in \mathbf{N}^d, \ \beta \in \mathbf{N}^m,$$ as new variables, together will all {\it algebraic} dependence relations
among the
functions \\
$x_1,...,x_d,
h_1(x),...,h_m(x)$. For instance it may
happen that $h_i(x)$ is a characteristic function of a Borel set,
in which case $h_i^2 = h_i$, or that $h(x)$ is an $m$-tuple of
algebraic functions, in which case a polynomial dependence
$P(x,h(x)) = 0$ holds. When no algebraic dependence exists
between the generators $x_1,...,x_d,
h_1(x),...,h_m(x),$ this relaxation method will treat them as independent variables.

More details and examples can be found in the recent preprint \cite{LP}.

\end{document}